An observer, whose eyes are 1.63 m above the ground, is standing 33.3 m away from a tree. The ground is level, and the tree is growing perpendicular to it. The observer's line of sight with the treetop makes an angle of 19.3° above the horizontal. How tall is the tree
1
Expert's answer
2017-09-15T15:36:07-0400
h=h_0+l*tanφ=1.63+33.3*tan⁡(19.3°)=13.29 m